# Attention-based Class-Conditioned Alignment for Multi-Source Domain Adaptive Object Detection

### v1

@@ Abstract
Domain adaptation methods for object detection (OD) strive to mitigate the impact of distribution shifts by promoting feature alignment across source and target domains. Multi-source domain adaptation (MSDA) allows leveraging multiple annotated source datasets, and unlabeled target data to improve the accuracy and robustness of the detection model. Most state-of-the-art MSDA methods for OD perform feature alignment in a class-agnostic manner. This is challenging since the objects have unique modal information due to variations in object appearance across domains. A recent prototype-based approach proposed a class-wise alignment, yet it suffers from error accumulation due to noisy pseudo-labels which can negatively affect adaptation with imbalanced data. To overcome these limitations, we propose an attention-based class-conditioned alignment scheme for MSDA that aligns instances of each object category across domains. In particular, an attention module coupled with an adversarial domain classifier allows learning domain-invariant and class-specific instance representations. Experimental results on multiple benchmarking MSDA datasets indicate that our method outperforms the state-of-the-art methods and is robust to class imbalance. Our code is available here.
Keywords:
Object Detection Multi-Source Domain Adaptation
@@ 1 Introduction
Given the recent advancements in deep learning, deep object detectors have shown impressive results on many benchmark problems ranging from closed-world fixed class settings to open-world scenarios [5, 41, 30, 19]. However, their performance typically degrades when there is a domain shift between the training (source) and test (target) data distributions. To address this challenge, several unsupervised domain adaptation (UDA) methods [16, 25, 4, 10] have been proposed to improve OD performance in the presence of a domain shift.
@@
The conventional UDA methods assume that the source data comes only from a single distribution. In real-world scenarios, the source dataset often comes from multiple distributions with considerable domain shifts due to weather, geographic location, and illumination. This setting is referred to as Multi-Source Domain Adaptation (MSDA) for OD. MSDA is challenging as it requires dealing with discrepancies among sources along with source and target discrepancies. However, having source data from multiple distributions results in enhanced adaptation, increased generalization, and improved robustness against domain shifts[21, 38]. Inspired by the UDA techniques, the typical approach for MSDA involves learning task-specific knowledge from the labeled source datasets, simultaneously trying to make the feature-space domain-invariant that generalizes well to the target domain. Several MSDA methods have been proposed for classification [28, 39, 1], yet MSDA for OD still remains relatively unexplored. This is because in the presence of multiple distribution shifts establishing the trade-off between learning domain-invariant features while preserving task-specific knowledge [15] is challenging when performed both at the image-level and instance level (for OD) compared to only at the image-level (for classification).
(a) No class-wise as in [35, 32]
(b) Prototype based [3]
(c) Ours
@@
Figure 1: Comparision of alignment of different MSDA techniques. (a) DMSN and TRKP implement pairwise-alignment of source-target pair without considering class-wise alignment. (b) Prototype based alignment (PMT) learns specific prototypes to represent each class and domain. Then, the same class prototype for different domains are merged in class-conditioned (domain invariant) prototypes. (c) In contrast, our ACIA learns instance-level domain invariant features by conditioning an attention module to attend a given class.
(a) No class-wise as in [35, 32]
(b) Prototype based [3]
(c) Ours
@@
Figure 2: T-SNE projection of the class distributions on BDD100K of source and target data. (a) Projection without class-specific alignment as in DMSN and TRKP. The classes are not aligned between the two domains. (b) Prototype-based alignment (PMT) uses prototypes for class alignment. Most classes are well aligned but the bike, as it is underrepresented, is not aligned. (c) Our ACIA uses an attention-based adversarial class alignment and manages to align all classes.
Only three methods have been proposed for MSDA of OD - Divide and Merge Spindle Network (DMSN) [35], Target Relevant Knowledge Preservation (TRKP) [32], and Prototype Mean Teacher (PMT) [3]. DMSN and TRKP focus on pairwise alignment between each source and target pairs as shown in Fig. 0(a). This ensures the alignment of each source domain with the target domain. However in a MSDA setting, there are considerable variations in object appearances across domains, so it’s also essential to focus on the class alignment [34]. Else it would lead to class misalignment across domains, which can be observed in Fig. 1(a) [^1] As both methods didn’t provide code or trained model, we used our method without instance-level alignment to generate the TSNE..
@@
Later, PMT proposed a method that uses prototypes for class-conditioned alignment. They use class-specific local and global prototypes to explicitly ensure within-domain and across-domain alignment, as summarized in Fig. 0(b). However, this method has a couple of critical limitations. Firstly, they make a strong preassumption that a single prototype can represent the true distribution of a class. But, generally, a class representation is not isotropic, which can be seen in the Fig. 1(b). Secondly, they rely on noisy pseudo-labelled target domains to learn the prototypes, resulting in class-misalignment due to error accumulation from the pseudo-labels [6] (we showed that in Tab. 5(right)). This issue is particularly problematic for underrepresented classes, which have fewer samples and can lead to misaligned class representations (emperical result in Tab. 6). This misalignment can be observed in Fig. 1(b), where the Bike class (red) is not aligned across the source and target domains. In this scenario, if we calculate the global prototype, it would lie somewhere in the middle of these two distributions, which does not accurately represent the actual class distribution. Consequently, this results in the degradation of the detector’s performance.
@@
This highlights the need for a better class-conditioned instance alignment mechanism in MSDA for OD. In this paper, we propose an Attention-based Class-wise Instance Aligner (ACIA) for MSDA, aiming at efficiently aligning class instances across domains, without relying on prototypical representation. The alignment process of ACIA is illustrated in Fig. 0(c) where each object category is aligned together across domains. In ACIA, the instances of each ROI-pooled features are infused with the class information by an attention operation. The class information is learned by using a label embedding. To compute the attention output, the query and key values are obtained from ROI-Pooled features, and the key is obtained from label embeddings. The output of the attention is used by adversarial domain discriminatior. The gradients are reversed from the output of the instance-level feature extractor forcing those features to be domain-invariant and class-wise aligned at the same time. Our method relies on just a single class embedding to store relative class information for each object category across domains. This allows a simple yet efficient multi-source class-conditional alignment. The benefits of our class-conditioned aligner can also be observed from Fig. 2, where our ACIA provides better class alignment across domains and robustness to imbalanced data.
@@
(1) An MSDA method called ACIA is proposed that integrates a conceptually simple class-wise instance alignment scheme. Instance alignment is implemented with an attention mechanism where instances from all domains are aligned to the corresponding class by the attention component.
@@
(3) Extensive experiments are conducted on three benchmark MSDA settings to validate the proposed ACIA alignment and ablation of the components. ACIA achieves state-of-the-art performance on all three benchmarks while remaining robust when adapting to imbalanced data. It is the simplest class-conditioned alignment process among the MSDA methods for OD.
2 Literature Review
@@
1. Unsupervised Domain Adaptation. UDA methods aim to learn a model that performs well on the target domain by combining the information from the labeled source data and unlabeled target data. A common approach for UDA is using an adversarial discriminator to generate domain-invariant representation. [8] integrated a GRL-based (reverses the gradient during backpropagation) discriminator at the image and instance level to align the features across domains. Later, [24] showed that strong and weak alignment of high-level and low-level features can improve the performance in UDA. [42] emphasizes the fact that in OD not all parts of the image are equally important. So, they focus on aligning the parts of the images that are more important using clustering and adversarial learning. [40] argues that aligning foreground regions is more important than background regions. They used the output of RPN as an attention map and prototypes for instance-level alignment. [14, 7, 23] tried to mitigate domain shift at the image level. They used Cycle-GAN to convert the source domain to the target domain and minimize the domain shift. These works focus on a single source, however, in this paper, we focus on the MSDA setting where the source data itself has multiple distribution shifts.
@@
2. Multi-Source Domain Adaptation. MSDA methods seek to adapt a model by using samples from multiple labeled source datasets and unlabeled target data. Only three works have been proposed to solve the problem of MSDA in OD - DMSN [35], TRKP[32], and PMT[3]. [35] learns pseudo-subnets for each domain and uses adversarial alignment to align each source subnet with the target. The weight of the target subnet is obtained as the weighted exponential average of the student subnets. [32] proposed an architecture with one detection head for each source domain. These detection heads are trained by using adversarial disentanglement for cross-domain alignment. They also introduced a mining strategy to weigh each source image based on its relevance to the target domain. Later [3] propose to use class and domain-specific prototypes as a representative of each class per domain. These prototypes are aligned by using contrastive loss to achieve class-conditioned adaptation. Out of the three methods only [3] focuses on class-wise alignment. However, they rely on the assumption that class representations are isotropic and also they suffer from pseudo-label error accumulation[6]. In contrast, our ACIA uses an attention-based adversarial aligner that is immune to pseudo-label error accumulation.
@@
3. Instance-Level Alignment in Domain Adaptation. The instance-level alignment is applied in a DA setting to ensure domain-invariant features at the instance level. [8] uses an adversarial domain classifier at the instance level similar to the image-level discriminator. [13, 33] improved on the work of [8] by giving weight to each ROI-pooled feature during instance-level alignment. This ensures that more importance is given to hard-to-classify features. [42] uses a COGAN to regenerate the ROI-pooled features and later align them using adversarial training. These works align the instance-level features in a class-agnostic way. [29] emphasizes the importance of object category-wise alignment. For this, they used several memory bank prototypes to store class information across domains and they are aligned by using class-specific discriminators. [37] learns a prototype for foreground and background across source and target domain. These prototypes are aligned to produce better predictions on the target domain. [34] learns graph-induced prototypes and proposes a loss to align them. [40] used prototypes and proposed to do instance-level alignment in a class-wise fashion using a prototype-based method. Nevertheless, all these methods rely on learning prototypes for instance-level class alignment in OD, based on the assumption that class representation is isotropic and used pseudo-labels for alignment.
@@ 3 Proposed Method
The MSDA setting consists of $N$ source domains $S_{1},S_{2},…,S_{N}$ and the target domain $T$. Similar to UDA setting, the object categories across all domains are fixed to a pre-defined set of categories and the number of object categories is $K$. Each source domain $S_{j}$ consists of images and corresponding labels, denoted as $S_{j}=\{(x^{j}_{i},y^{j}_{i})\}^{M_{j}}_{i=1}$, where $j=1,2,...,N$, and $M_{j}$ is the number of images in the source domain $j$. Here, $x^{j}_{i}$ is the $i$th image from source domain $S_{j}$ and $y^{j}_{i}$ is the set of object labels and corresponding bounding boxes. The target domain $T$ only has images, with no bounding boxes and class labels information. It can be represented as $T={\{x_{i}^{T}\}^{M_{T}}_{i=1}}$, where $M_{T}$ is the number of images in the target domain and $x_{i}^{T}$ is the $i$th image from the target domain.
The training framework of our proposed ACIA is shown in Fig. 3. Our model is based on the Mean-Teacher (MT) framework[10]. The components responsible for image-level and instance-level alignment are integrated into the MT framework. They are the multi-source discriminator and attention-based class-conditioned instance-level discriminator. The MT training process is performed in two stages. First, the student model is trained in a supervised manner on all source domain data, typically called the burn-in stage in the MT training process [10]. After this step, the teacher model weights are initialized as an exact copy of the student model weights. Then, mutual learning of teacher-student begins on both the labeled source datasets and unlabeled target data. In the case of source datasets, the same supervised training is continued. For the target dataset images, two augmented versions are created for each image, called the weak and strong augmentation. The teacher model receives the weak augmented version and it computes the pseudo-labels. The student model receives the strong augmented version and computes its loss with the pseudo labels provided by the teacher as ground truth. Along with that, the multi-source discriminator and class-conditioned instance-level discriminator are trained for domain-invariant feature generation and class-wise alignment respectively. The weight of the teacher model is the exponential moving Average(EMA) of the weight of the student model. At the time of inference, only the teacher model is used. The rest of the section describes all the loss functions.
@@
1. Source Data Training The source data from all the domains have annotations available. Thus the standard supervised training is applied for the source datasets where the student network weights are updated by backpropagating the detection loss:
@@
where $\mathcal{L}_{\mbox{reg}}$ is the smooth-L1 loss [22] used for the bounding box regression and $\mathcal{L}_{\mbox{cls}}$ is the focal loss [17] used for bounding box classification. The source datasets are used in both the training stages, ie, the burn-in and teacher-student mutual learning.
@@
2. Target Data Training Target datasets are used in the training process at the teacher-student mutual learning stage. For every image from the target domain, two augmented versions are generated - strong and weak augmentation. For the strong augmentation, we used color jittering, grayscale, Gaussian blur, and cutout patches. For weak augmentation, image rescaling and horizontal flips are used. The weak augmented image is fed to the teacher model and the output is post-processed to generate pseudo-labels $\widetilde{y}$. The post-processing applied is the confidence thresholding [10]. Let the output of the teacher network for an image $x_{i}$ be $\hat{y}_{i}=\{c_{i}^{j},p_{i}^{j},b_{i}^{j}\}_{j=1}^{D}$ where $c_{i}^{j},p_{i}^{j},b_{i}^{j}$ are predicted class, probability and box respectively and $D$ is the number of output detections. Then $\widetilde{y}_{i}=\{c_{i}^{j},b_{i}^{j}|p_{i}^{j}>\theta\}$ where $\theta$ is a threshold used for filtering noisy pseudo-labels. The student network then receives the strongly augmented version and computes its loss. The loss function used is the same as eqn. 1, except the pseudo-labels are employed:
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x_{i},\widetilde{y}_{i})+\mathcal{L}_{\mbox{reg}}(x_{i},\widetilde{y}_{i})$ (2)
Once the student weights are updated via backpropagation, the teacher weights are updated as $w_{t}=\alpha w_{t}+(1-\alpha)w_{s}$ where $w_{t},w_{s}$ are the weights of teacher and student networks respectively and $\alpha$ is a smoothing factor.
@@
3. Image-Level Alignment In the domain adaptation problem, the task-specific knowledge is learned from the source domain. But, for the model to perform well on the target domain, we have to learn a representation that is also domain-invariant. In a UDA setting, this is achieved by attaching a domain discriminator (binary classifier) to the output of the backbone using a gradient-reversal layer (GRL) [16]. This results in a min-max optimization game between the backbone and the domain-discriminator, leading to domain-invariant feature representation. Unlike UDA, in MSDA setting the source data comes from various distributions, so we used a multi-class discriminator instead of a binary discriminator. In our work, we used a CNN-based classifier that outputs a domain prediction map from the last feature map of the backbone. The cross-entropy loss is then applied to all pixels of the prediction map and the results are pooled using global average pooling. The training loss of the image-level discriminator can be represented as:
$\mathcal{L}_{\mbox{dis}}=-\frac{1}{WH}\sum_{j=1}^{N+1}\sum_{i=1}^{M_{j}}\sum_{h=1,w=1}^{H,W}d_{i}^{j}\log(D(F(x_{i}^{j})_{wh})$ (3)
@@
Here, W and H are the width and height of the prediction map. $D(F(x_{i}^{j})_{wh}$ is the output of the multi-class discriminator for pixel $w,h$ of the prediction map. D and F represent the discriminator and Feature extractor (the backbone network in this case) respectively. $d_{i}\in\{0,1\}^{N+1}$ whose value is one for the correct domain of the $i$th image and zero everywhere else.
@@
4. Instance-level Alignment The instance-level alignment is performed in a class-wise fashion in our method. Class-wise alignment helps to align features of objects from the same class regardless of the difference in their visual attributes (such as appearance, scale, orientation, and other characteristics of objects) both within and across domains. Thus even in the presence of class imbalance, underrepresented classes will be focused on ensuring the alignment within and across domains. As shown in fig 2, without class-wise alignment, the separation of class features of the same category across domains is high. PMT [3] is the first MSDA approach in OD that implements class-wise alignment using global and local prototypes. The local prototypes ensured within domain alignment whereas the global prototypes ensured across domain alignment. In contrast to using prototypes for each object category and each domain, we use an efficient attention-based module to ensure class-wise instance alignment.
@@
The instance-level alignment in OD is performed using the region features from each instance. Different from image-level alignment, we have features for a single object category at the instance level. Following the previous approaches, an adversarial discriminator on RoI-Pooled features can be used to perform the instance-level feature alignment. But this alignment process is class-agnostic. If we can perform this adversarial learning in a class-wise fashion, the resulting feature alignment will be class-conditioned. The class information is available for the source domain from the annotations. For the target domain, we can use the filtered pseudo-labels generated by the teacher model. However, the pseudo-labels are noisy, even after filtering. Using them would lead to accumulation error during training from the noises in pseudo-labels [6]. We empirically also found that using the pseudo-labels in class-conditioned alignment harms the model’s performance (see Tab. 5).
@@
To use class information during the alignment process, we create learnable embeddings for each class label. These embeddings can be coupled with region embeddings from the RoI Pool using simple operations like concatenation or multiplication. It is observed that the attention mechanisms are effective at finding relations between two vectors by dynamically focusing on different parts of the sequences [2, 20]. So, we decided to employ an attention-based alignment operation (the results from empirical comparison are available in table 5). For this, two linear projections of the ROI-Pooled features and a linear projection from the label embedding are created. The projections from region embedding are used as query($Q$) and value($V$) respectively, while the projection from label embedding is used as key($K$) for the attention block (see Fig. 3). The attention block uses a multi-head attention as proposed in the original paper [27] which can be expressed as:

### v2

# Attention-based Class-Conditioned Alignment for Multi-Source Domain Adaptation of Object Detectors

@@ Abstract
Domain adaptation methods for object detection (OD) strive to mitigate the impact of distribution shifts by promoting feature alignment across source and target domains. Multi-source domain adaptation (MSDA) allows leveraging multiple annotated source datasets and unlabeled target data to improve the accuracy and robustness of the detection model. Most state-of-the-art MSDA methods for OD perform feature alignment in a class-agnostic manner. This is challenging since the objects have unique modal information due to variations in object appearance across domains. A recent prototype-based approach proposed a class-wise alignment, yet it suffers from error accumulation due to noisy pseudo-labels that can negatively affect adaptation with imbalanced data. To overcome these limitations, we propose an attention-based class-conditioned alignment method for MSDA that aligns instances of each object category across domains. In particular, an attention module coupled with an adversarial domain classifier allows learning domain-invariant and class-specific instance representations. Experimental results on multiple benchmarking MSDA datasets indicate that our method outperforms the state-of-the-art methods and is robust to class imbalance using a conceptually simple class-conditioning method. Our code is available here.
@@ 1 Introduction
Given the recent advancements in deep learning, object detectors have demonstrated impressive results on many benchmark problems ranging from closed-world fixed class settings to open-world scenarios [5, 42, 31, 20]. However, their performance typically degrades when there is a domain shift between the training (source) and test (target) data distributions. To address this challenge, several unsupervised domain adaptation (UDA) methods [17, 26, 4, 10] have been proposed to improve OD performance in the presence of a domain shift. The conventional UDA methods assume that the source data comes from a single distribution. In real-world scenarios, multiple source datasets may be available for adaptation with considerable domain shifts among these datasets due to variations in capture conditions, like weather, geographic location, and illumination. This setting is referred to as Multi-Source Domain Adaptation (MSDA). MSDA is challenging as it requires addressing discrepancies among different sources, along with sources and target discrepancies. However, leveraging source data from multiple distributions results in enhanced adaptation, increased generalization, and improved robustness against domain shifts[22, 39].
@@
Inspired by UDA methods, the typical approach for MSDA involves learning task-specific knowledge from the labeled source datasets, simultaneously trying to learn a domain-invariant feature representation that generalizes well to the target domain. Several MSDA methods have been proposed for classification [29, 40, 1], yet MSDA for OD remains relatively unexplored. This is due to the challenges of establishing a good trade-off between learning domain-invariant features while preserving task-specific knowledge [16] when performed both at the image-level and instance-level (for OD), compared to only at the image-level (for classification).
(a) Methods w/o class alignment [36, 33].
(b) Prototype-based method [3].
(c) Our ACIA method.
@@
Figure 1: Comparison of alignment strategies of different MSDA methods. (a) DMSN and TRKP implement pairwise alignment of source-target pair without considering class-wise alignment. (b) PMT learns specific prototypes to represent each class and domain. Then, the same class prototypes from different domains are merged into class-conditioned domain-invariant prototypes. (c) In contrast, our ACIA learns instance-level domain invariant features by conditioning an attention module to attend a given class.
(a) Methods w/o class alignment [36, 33].
(b) Prototype-based method [3].
(c) Our ACIA method.
@@
Figure 2: T-SNE projection of the class distributions of source data (BDD100K Daytime) and target data (BDD100K Dusk/Dawn). (a) Projection without class-specific alignment as in DMSN and TRKP. The classes are not aligned between the two domains. (b) PMT relies on prototypes for class alignment. Most classes are well aligned except for the bike since it is underrepresented. (c) Our ACIA uses an attention-based adversarial class alignment and manages to align all classes.
@@
Only three methods have been proposed for MSDA of OD - Divide and Merge Spindle Network (DMSN) [36], Target Relevant Knowledge Preservation (TRKP) [33], and Prototype Mean Teacher (PMT) [3]. DMSN [36] and TRKP [33] focus on pairwise alignment between each source and target pairs as shown in Fig. 1(a). It ensures the alignment of each source domain with the target domain. However, in an MSDA setting, there are considerable variations in object appearances across domains, so it’s also essential to focus on the class alignment [35, 3]. Otherwise, this would lead to class misalignment across domains, as can be observed in Fig. 2(a) [^1] Since both methods do not provide their code or trained model, we used our method without instance-level alignment to generate t-SNE plots.. Later, PMT [3] was proposed that uses prototypes for class-conditioned alignment. It relies on class-specific local and global prototypes to explicitly ensure within-domain and across-domain alignment, as summarized in Fig. 1(b). However, this method has some important limitations. Firstly, it assumes that a single prototype can represent the true distribution of a class. But, generally, a class representation is not isotropic, which can be observed in the Fig. 2(b). Secondly, it relies on noisy pseudo-labelled target domains to learn the prototypes, resulting in class-misalignment due to error accumulation from the pseudo-labels [6] (see Tab. 6 (right)). This issue is particularly problematic for underrepresented classes, which have fewer samples and can lead to misaligned class representations (see the empirical result in Tab. 11). This misalignment can be observed in Fig. 2(b), where the Bike class (red) is not aligned across the source and target domains. In this scenario, the global prototype would lie somewhere between these two distributions, which does not accurately represent the class distribution. This results in a degradation of detector performance.
@@
The above issues highlight the need for MSDA methods for OD with better class-conditioned instance alignment mechanisms. In this paper, an Attention-based Class-wise Instance Aligner (ACIA) for MSDA is proposed to efficiently align class instances across domains without relying on prototypical representation. The alignment process of ACIA is illustrated in Fig. 1(c), where each object categories are aligned across domains. The instances of each ROI-pooled feature are fused with the class information using an attention operation. The class information is learned by using a label embedding. To compute the attention output, the query and the value are obtained from ROI-pooled features, and the key is obtained from label embedding. The output of the attention module is fed to a domain discriminator through the Gradient Reversal Layer (GRL) [11]. This reverses the gradient of the output of the instance-level features, forcing them to be domain-invariant and class-wise aligned simultaneously. ACIA relies on just a single class embedding to store relative class information for each object category across domains. This allows for a simple yet efficient multi-source class-conditional alignment. The benefits of our class-conditioned aligner can also be observed in Fig. 2, where ACIA provides better class alignment across domains and robustness to imbalanced data.
@@
An MSDA method for OD called ACIA is proposed that integrates a conceptually simple class-wise instance alignment scheme. It implements instance alignment with an attention mechanism, where instances from all domains are aligned to the corresponding class.
@@
(3) An extensive set of experiments is conducted on three benchmark MSDA settings to validate the proposed ACIA alignment and characterize its key components. ACIA achieves state-of-the-art performance on all three benchmarks while remaining robust when adapting to imbalanced data. It is the most cost-effective MSDA method for OD.
2 Related Work
@@
(a) Unsupervised Domain Adaptation. UDA methods aim to learn a model that performs well on the target domain by combining the information from labeled source data and unlabeled target data. A common approach for UDA is using an adversarial discriminator to generate domain-invariant representation. [8] integrated a discriminator based on GRL (reverses the gradient during backpropagation) at the image and instance level to align the features across domains. Later, [25] showed that strong and weak alignment of high- and low-level features can improve the UDA performance. [43] focuses on aligning the parts of the images that are more important using clustering and adversarial learning. [41] argues that aligning foreground regions is more important than background regions. They used the output of RPN as an attention map and prototypes for instance-level alignment. [15, 7, 24] tried to mitigate domain shift at the input level. They used Cycle-GAN to convert the source to the target domain and minimize the domain shift. Although these works focus on a single source, they have provided important mechanisms to develop MSDA methods where the source data has multiple distribution shifts.
@@
(b) Multi-Source Domain Adaptation. MSDA methods seek to adapt a model by using samples from multiple labeled source datasets and unlabeled target data. Only three works have been proposed to solve the problem of MSDA in OD – DMSN [36], TRKP [33], and PMT [3]. [36] learns pseudo-subnets for each domain and uses adversarial alignment to align each source subnet with the target. The weight of the target subnet is obtained as the weighted exponential average of the student subnets. [33] proposed an architecture with one detection head for each source domain. These detection heads are trained by using adversarial disentanglement for cross-domain alignment. A mining strategy is also used to weigh each source image based on its relevance to the target domain. Later [3] propose to use class and domain-specific prototypes, which are aligned by using contrastive loss to achieve class-conditioned adaptation. Out of the three methods, only [3] focuses on class-wise alignment. However, they rely on the assumption that class representations are isotropic, and they suffer from pseudo-label error accumulation [6], especially when the dataset is imbalanced. In contrast, our ACIA relies on an attention-based adversarial aligner that is immune to pseudo-label error accumulation and robust to imbalanced datasets.
@@
(c) Instance-Level Alignment in Domain Adaptation. Instance-level alignment is applied in the UDA setting to ensure domain-invariant features at the instance level. [8] uses an adversarial domain classifier at the instance level similar to the image-level discriminator. [14, 34] improved on the work of [8] by giving weight to each ROI-pooled feature ensuring that more importance is given to hard-to-classify features. [43] uses a COGAN to regenerate the ROI-pooled features and later align them using adversarial training. These works align the instance-level features in a class-agnostic way. [30] uses several memory bank prototypes to store class information across domains and align them using class-specific discriminators. [38] aligns prototypes for foreground and background across source and target domains to produce better predictions on the target domain. [35] learns graph-induced prototypes and proposes a loss to align them. [41] used prototypes and proposed to do instance-level alignment in a class-wise fashion using a prototype-based method. Nevertheless, all these methods rely on learning prototypes for instance-level alignment in OD, based on the assumption that class representation is isotropic and uses pseudo-labels for alignment.
@@ 3 Proposed Method
The MSDA setting consists of $N$ source domains $S_{1},S_{2},…,S_{N}$ and a target domain $T$. Similar to the UDA setting, the object categories across all domains are fixed to a pre-defined set of $K$ categories. Each source domain $S_{j}$ consists of images and corresponding labels, denoted as $S_{j}=\{(x^{j}_{i},y^{j}_{i})\}^{M_{j}}_{i=1}$, where $j=1,2,...,N$, and $M_{j}$ is the number of images in the source domain $j$. Here, $x^{j}_{i}$ is the $i$th image from source domain $S_{j}$, and $y^{j}_{i}$ is the set of object labels and the corresponding bounding boxes (bboxes). The target domain $T$ only has images with no annotations. It can be represented as $T={\{x_{i}^{T}\}^{M_{T}}_{i=1}}$, where $M_{T}$ is the number of images in the target domain and $x_{i}^{T}$ is the $i$th image from the target domain.
The training framework of our proposed ACIA is shown in Fig. 3. Our model is based on the Mean-Teacher (MT) framework [10]. The components responsible for image-level and instance-level alignment are integrated into the MT framework. They are the multi-source discriminator and attention-based class-conditioned instance-level discriminator. The MT training process is comprised of two stages. First, the student model is trained in a supervised manner on all source domain data, typically called the burn-in stage. After this step, the teacher model weights are initialized as an exact copy of the student model weights. Then, mutual learning of teacher-student begins on both the labeled source datasets and unlabeled target data. In the case of source datasets, the same supervised training is continued. For the target dataset images, pseudo labels generated by the teacher model are used for training. Meanwhile, the multi-source discriminator and class-conditioned instance-level discriminator are trained for domain-invariant feature generation and class-wise alignment, respectively. The weights of the teacher model are the exponential moving average (EMA) of the student model weights. At inference time, only the teacher model is employed. The rest of the section describes all the loss functions for the proposed ACIA framework.
@@
(a) Source Data Training. Data from all the source domains are annotated. Thus, the standard supervised training is applied on the source datasets where the student network weights are updated through backpropagation of the detection loss:
@@
where $\mathcal{L}_{\mbox{reg}}$ is the smooth-L1 loss [23] used for the bboxes regression and $\mathcal{L}_{\mbox{cls}}$ is the focal loss [18] used for bboxes classification.
@@
(b) Target Data Training. The target dataset is used at the teacher-student mutual learning stage of the training process. For every image from the target domain, two augmented versions are generated – with strong and weak augmentation. For the strong augmentation, color jittering, grayscale, Gaussian blur, and cutout patches are used. For weak augmentation, image rescaling, and horizontal flips are used. The weak augmented image is fed to the teacher model, and the output is filtered by confidence thresholding [10] to generate pseudo-labels $\widetilde{y}$. The student network then receives the strongly augmented version and its loss is computed using the the same loss function as Eq. 1:
$\mathcal{L}_{\mbox{unsup}}=\sum_{i=1}^{M_{T}}\mathcal{L}_{\mbox{cls}}(x^{T}_{i},\widetilde{y}_{i})+\mathcal{L}_{\mbox{reg}}(x^{T}_{i},\widetilde{y}_{i})$ (2)
@@
(c) Image-Level Alignment. In the UDA setting, task-specific knowledge is learned from the source domain. However, for the model to perform well on the target domain, a domain-invariant representation must also be learned. This is achieved by attaching a domain discriminator (binary source-target classifier) to the output of the backbone that uses a GRL [17]. This results in a min-max optimization game between the backbone and the domain-discriminator, leading to domain-invariant feature representation. In the MSDA setting, the source data comes from multiple distributions, and therefore, a multi-class discriminator is required. In our work, we used a CNN-based classifier that outputs a domain prediction map from the last feature map of the backbone. The cross-entropy loss is then applied to all pixels of the prediction map, and the results are pooled using global average pooling. The training loss of the image-level discriminator can be expressed as:
$\mathcal{L}_{\mbox{dis}}=-\frac{1}{WH}\sum_{j=1}^{N+1}\sum_{i=1}^{M_{j}}\sum_{h=1,w=1}^{H,W}d_{i}^{j}\log(D_{M}(F(x_{i}^{j})_{wh})$ (3)
@@
where W and H are the width and height of the prediction map. $D_{M}$ and $F$ represent the multi-class discriminator and feature extractor (backbone network), respectively. $D_{M}(F(x_{i}^{j})_{wh}$ is the output of the multi-class discriminator for pixel $w,h$ of the prediction map. $d_{i}^{j}$ is a ${N+1}$ sized one-hot vector, where the position corresponding to the domain of the $i$th image is set to 1 and 0 everywhere else.
@@
(d) Instance-level Alignment. In our method, the instance-level alignment is performed in a class-wise manner. Each object category has variations in visual attributes (such as appearance, scale, orientation, and other characteristics) across domains. If the variation is significant, it is challenging for the detector to align features across domains, leading to sub-optimal performance [35, 3]. Furthermore, class-wise alignment helps with underrepresented classes by ensuring that the features of these classes are explicitly aligned across domains, regardless of their lower frequency in the data. As shown in Fig. 2, without class-wise alignment, the separation of class features of the same category across domains is high. PMT [3] is the first MSDA approach in OD that implements class-wise alignment using global and local prototypes. The local prototypes ensured within-domain alignment, whereas the global prototypes ensured across-domain alignment. In contrast to using prototypes for each object category and each domain, we use an efficient attention-based module to ensure class-wise instance alignment.
@@
The instance-level alignment in OD is performed using the RoI-Pooled features from each instance. Unlike image-level alignment, each feature at the instance level represents only a single object category. Following previous approaches, an adversarial discriminator on RoI-Pooled features can perform the instance-level feature alignment. But this alignment process is class-agnostic. If we can perform this adversarial learning class-wise, the resulting feature alignment will be class-conditioned. For the source data, class information is available from their annotations. For the target domain, we can use the filtered pseudo-labels generated by the teacher model. However, the pseudo-labels are noisy, even after filtering. Using them would lead to an accumulation of errors during training from the noise in pseudo-labels [6]. In addition, we empirically found that using the pseudo-labels in class-conditioned alignment harms model performance (see Tab. 6).
@@
To use class information during the alignment process, we create learnable embeddings for each object category. Each embedding is responsible for learning class information corresponding to an individual object category. The class information from embeddings can be coupled with region embeddings from the RoI-Pool (resulting in class-aware region features) using simple operations like concatenation or multiplication. It is observed that the attention mechanisms are effective at finding relations between two vectors by dynamically focusing on different parts of the sequences [2, 21]. In our ACIA, we employ an attention-based alignment operation. (The results from the empirical comparison are available in Tab. 6.) For this, two linear projections of the ROI-Pooled features and a linear projection from the label embedding are created. The projections from ROI-Pooled features are used as query ($Q$) and value ($V$), while the projection from label embedding is used as key ($K$) for the attention block (see Fig. 3). The attention block uses multi-head attention as proposed in the original paper [28] which can be expressed as: